# Synthetic POMDPs to Challenge Memory-Augmented RL: Memory Demand Structure Modeling

### v1

@@ 1 Introduction
In reinforcement learning (RL), Partially Observable Markov Decision Process (POMDP) serves as a canonical framework for modeling real-world decision-making, where agents must infer hidden Markov states from past observations. Memory-augmented RL algorithms, equipped with memory models such as LSTM (Hochreiter and Schmidhuber 1997), have emerged as a key solution to tackle POMDPs. By introducing memory, these algorithms can summarize past observations, thereby restoring policy convergence guarantees for POMDPs (Sutton and Barto 2018).
Prior research has devised a diverse range of benchmarks that evaluate RL algorithms from multiple perspectives (Bellemare et al. 2013; Brockman et al. 2016; Osband et al. 2019; Küttler et al. 2020; Cobbe et al. 2020; Raffin et al. 2021; Hafner 2021; Morad et al. 2023; Huang et al. 2024; Tao et al. 2025). Judging from the proportion of POMDP environments in these benchmarks, the modeling and solving of POMDPs with RL has gained increasing academic attention. Notably, these benchmarks encompass classic control problems, game-based tasks, and other scenarios, some of which include synthetic environments.
Synthetic environments, designed for targeted testing, include scenarios such as applying masks/noise to Markov states to assess handling of partial observability, designing sparse reward signals to evaluate exploration efficiency, or constructing memory tasks to examine memory retrieval capabilities. While environments abstracted from real-world problems are sufficiently complex to evaluate the holistic performance of memory-augmented RL effectively, synthetic environments remain indispensable. Specifically, their inherently simple and controllable mechanisms grant researchers greater freedom to engineer task dynamics, unencumbered by the intertwined complexities of real-world problems. This controllability facilitates targeted and independent assessment of specific performance dimensions. Given the diverse complexities of real-world problems and the lack of consensus on a universal memory model for RL, these targeted assessments retain critical academic value.
While adopting synthetic POMDP environments has become a common practice in existing benchmarks, critical limitations persist in this field. First, the absence of a theoretical framework to characterize the difficulty of POMDPs for memory models leaves the design of new environments reliant on intuition and experience. This raises questions about whether synthetic environments in current benchmarks adequately cover diverse Memory Demand Structures (MDS) and difficulty levels. Second, there is a lack of universal methodology for constructing synthetic POMDPs to test memory-augmented RL, alongside insufficient fine-grained control over environmental characteristics.
Specifically, two primary approaches currently dominate efforts to manipulate synthetic POMDP difficulty: 1. Controlling observable state information via masking (Raffin et al. 2021; Morad et al. 2023; Tao et al. 2025). 2. Modifying problem scale (e.g., steps, state/action space size) or adjusting randomness (Osband et al. 2019; Cobbe et al. 2020; Morad et al. 2023; Tao et al. 2025). The first approach, although intuitive and easy to implement, fails to capture the decision value of distinct information fields accurately. Consequently, the resulting difficulty levels are limited in number and unevenly distributed. The second method modifies a single problem into multiple difficulty levels, but performance degradation from increased difficulty cannot be attributed solely to memory model deficiencies; it may also stem from RL algorithm limitations. This is because such methods alter the underlying belief MDP (i.e., the MDP over belief states in POMDPs) (Kaelbling, Littman, and Cassandra 1998) while changing the MDS. The belief MDP measures a problem’s intrinsic difficulty for RL algorithms, excluding the difficulty of extracting trajectory information to construct belief states. To evaluate RL memory models, adjusting difficulty should ideally leave the underlying belief MDP unchanged. Our review, however, did not find work addressing this.
Our study aims to address the limitations mentioned above by proposing a principled framework for synthesizing POMDPs to evaluate memory models in RL algorithms. The main contributions of this work are as follows:
@@
for analyzing POMDP through MDS and transition invariance (stationarity and consistency). MDS captures the memory demands imposed by a trajectory, while consistency and stationarity capture the fitting requirements for memory models across the observational and temporal dimensions of trajectories.
@@
leveraging linear process dynamics, state aggregation, and reward redistribution, enabling the customization of POMDPs with specified memory challenges. Among them, linear process dynamics directly constructs POMDPs; state aggregation and reward redistribution build POMDPs from existing MDPs, with state aggregation preserving the belief MDP and reward redistribution preserving the optimal policy.
@@
validated through detailed experiments, demonstrating their effectiveness in distinguishing the capabilities of various memory models in RL.
Based on theory and empirical validation, our approach offers researchers actionable guidelines for designing POMDP environments, laying a rigorous foundation for evaluating memory models in RL.
@@ 8 Conclusion and Future Work
We developed an MDS-centered theory to analyze and construct POMDPs across difficulty levels. Integrating our transition invariance concept with tools like linear process dynamics, state aggregation, and reward redistribution, we built three series of increasingly difficult POMDP environments, where we tested RL with various memory models.
In our experiments, the simple Elman model outperformed more complex ones in most cases. Moreover, as MDS complexity increased, Elman’s performance degraded less severely than that of other models, indicating its advantage in capturing higher-order dependency information. This phenomenon helps explain Elman’s advantage reported in the POPGym paper (Morad et al. 2023).
We propose incorporating test series with increasing MDS complexity as part of the evaluation criteria for memory models in RL. Plotting performance curves such as Fig. 8 facilitates understanding the representational capacity of memory models in capturing higher-order dependencies within RL tasks. In future work, we will: 1. construct more diverse environment series, including cases where different trajectories exhibit varying MDS; 2. conduct additional experiments to guide RL tasks in selecting from a broader range of memory models.

### v2

@@ 1 Introduction
Within the domain of reinforcement learning (RL), the Partially Observable Markov Decision Process (POMDP) provides a standard formalism for simulating decision-making in practical settings, requiring agents to deduce latent system states from historical data. RL agents enhanced with memory components, for instance Long Short-Term Memory networks (Hochreiter and Schmidhuber 1997), represent a principal strategy for addressing POMDPs. The incorporation of memory allows these agents to integrate historical observations, thus re-establishing guarantees for policy convergence within POMDP contexts (Sutton and Barto 2018).
A variety of benchmarking suites have been developed in prior work to assess RL algorithms across numerous facets (Bellemare et al. 2013; Brockman et al. 2016; Osband et al. 2019; Küttler et al. 2020; Cobbe et al. 2020; Raffin et al. 2021; Hafner 2021; Morad et al. 2023; Huang et al. 2024; Tao et al. 2025). The significant presence of POMDP-based tasks within these benchmarks indicates a growing research focus on addressing POMDPs using RL techniques. These suites incorporate traditional control issues, challenges derived from games, and other settings, a number of which are artificially constructed.
Synthetic environments are crafted for focused evaluation, featuring designs like obscuring or adding noise to true states to test an agent’s ability to handle incomplete information, implementing sparse rewards to gauge exploration strategies, or creating tasks that specifically probe memory recall. Although environments derived from real-world applications possess inherent complexity suitable for testing the overall efficacy of memory-augmented RL, synthetic settings are still crucial. Their straightforward and adjustable nature offers researchers enhanced flexibility to shape task mechanics, free from the entangled intricacies found in real scenarios. This capacity for precise control allows for isolated and directed measurement of particular algorithmic capabilities. Owing to the varied nature of practical challenges and the absence of an agreed-upon standard memory architecture for RL, such focused evaluations maintain substantial importance.
Despite the widespread incorporation of artificial POMDP settings in contemporary benchmarks, significant shortcomings remain. Firstly, the lack of a conceptual structure to define the challenge a POMDP poses to memory models means that creating new environments still depends heavily on heuristic judgment. This casts doubt on whether existing synthetic tasks sufficiently represent a wide spectrum of Memory Demand Structures (MDS) and complexity gradients. Secondly, a systematic procedure for building synthetic POMDPs to test memory-augmented RL is missing, coupled with inadequate precision in tuning environmental attributes.
Presently, two main strategies are predominantly used to vary the difficulty of synthetic POMDPs: 1. Regulating the observable state information through masking techniques (Raffin et al. 2021; Morad et al. 2023; Tao et al. 2025). 2. Altering the problem scope (such as episode length, dimensionality of state/action spaces) or introducing stochasticity (Osband et al. 2019; Cobbe et al. 2020; Morad et al. 2023; Tao et al. 2025). The first strategy, while straightforward to apply, does not adequately reflect the relative importance of different information elements for decision-making. As a result, the difficulty levels produced are often few and not uniformly spaced. The second approach generates multiple difficulty tiers from a base problem, but any performance drop with increased difficulty cannot be uniquely ascribed to memory model shortcomings; limitations of the core RL algorithm might also be a factor. This is because these methods change the fundamental belief MDP (the MDP defined over belief states in a POMDP) (Kaelbling, Littman, and Cassandra 1998) concurrently with altering the MDS. The belief MDP quantifies a problem’s inherent difficulty for an RL agent, separate from the challenge of inferring belief states from trajectories. For a clean evaluation of RL memory models, difficulty adjustments should, ideally, keep the underlying belief MDP invariant. Our examination of the literature found no prior work successfully achieving this.
This research seeks to overcome the stated limitations by introducing a structured framework for generating POMDPs tailored to evaluating memory models in RL. The key contributions of this paper are:
@@
for analyzing POMDPs based on MDS and transition invariance (encompassing stationarity and consistency). MDS quantifies the memory requirements dictated by a trajectory, whereas consistency and stationarity describe the fitting demands placed on memory models across observational and temporal scales.
@@
that utilizes linear process dynamics, state aggregation, and reward redistribution, allowing for the creation of POMDPs with predefined memory challenges. Herein, linear process dynamics builds POMDPs directly; state aggregation and reward redistribution derive POMDPs from pre-existing MDPs, with state aggregation maintaining the original belief MDP and reward redistribution preserving the optimal policy.
@@
that are empirically validated, showing their utility in discriminating the performance of different memory models used in RL.
Grounded in theoretical insights and experimental proof, our proposed method supplies practical guidance for crafting POMDP environments, establishing a solid basis for the assessment of memory models in RL.
@@ 8 Conclusion and Future Work
We have established a theoretical framework centered on Memory Demand Structures (MDS) to systematically analyze and generate Partially Observable Markov Decision Processes (POMDPs) with scalable difficulty levels. By integrating the concept of transition invariance—encompassing stationarity and consistency—with methodological tools such as linear process dynamics, state aggregation, and reward redistribution, we constructed three distinct series of POMDP environments. Each series exhibits a progressively increasing level of complexity, providing a structured testbed for evaluating reinforcement learning (RL) agents equipped with diverse memory architectures.
Experimental results revealed a notable trend: the relatively simple Elman network (RNN) frequently outperformed more sophisticated memory models across a majority of the evaluated settings. Furthermore, as the MDS complexity escalated within the environment series, the performance degradation observed in the Elman model was significantly less pronounced compared to other models. This suggests that the Elman network possesses a superior capability in capturing and leveraging higher-order temporal dependencies present in the observation sequences. These findings offer a plausible explanation for the performance advantage of Elman-style networks previously noted in studies such as the POPGym benchmark (Morad et al. 2023).
Based on these insights, we advocate for the inclusion of progressively challenging MDS-based test series as a standard component in the evaluation protocol for memory-augmented RL algorithms. Visualizing performance through curves, such as those depicted in Fig. 8, offers an intuitive means to assess and compare the ability of different memory models to represent complex, long-range dependencies critical for solving RL tasks under partial observability.
Looking forward, our future work will focus on two main directions: 1. Developing a more extensive and varied collection of environment series. This includes designing scenarios where different trajectories within the same environment exhibit heterogeneous MDS patterns, thereby presenting a more nuanced and realistic challenge. 2. Undertaking a broader set of empirical studies aimed at providing practical guidance for model selection. The goal is to establish clearer correlations between specific task characteristics, encoded by MDS profiles, and the most suitable memory model from an expanded set of candidates, ultimately aiding practitioners in optimizing their RL agent designs.